OpenStudy (anonymous):

what is 1 1/2 % of 250?

OpenStudy (anonymous):

1 1/2 = 3/2 = 1.5 (15x25)/100

OpenStudy (anonymous):

erica, do you have problem with percentage?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

% is 100, if you find 1.5, convert it to 15 and 1000 if 1.25 ---> 125 and 10000